2) The residents of Vegopia spend all of their income on cauliflower, broccoli, and carrots. In 2016, they spend a total of $210 for 100 heads of cauliflower, $70 for 50 bunches of broccoli, and $45 for 450 carrots. In 2017, they spend a total of $225 for 90 heads of cauliflower, $112 for 80 bunches of broccoli, and $120 for 400 carrots.

Complete the following table by calculating the price of one unit of each vegetable in each year


Complete the following table by calculating the price of one unit of YearCauliflowerBroccoll 2016$ 2017$ Carrots Using 2016 as the base year, the CPI for 2016 is ? , and The inflation rate in 2017 is %using the CPI.

Using 2016 as the base year, the CPI for 2016 is ___ , and the CPI for 2017 is ___   
The inflation rate in 2017 is____% using the CPI.

media%2F177%2F177cc402-e8c7-4da9-a1cd-a9
media%2F1b9%2F1b99c4af-5340-4a8b-9a7d-6a
Answer #1

In year 2016,

Price of one unit of cauliflower = 210/100 = $2.1

Price of one unit of Broccoli = 70/50 = $1.4

Price of one unit of carrots = 45/450 = $.1

In year 2017,

Price of one unit of cauliflower = 225/90 = $2.5

Price of one unit of Broccoli = 112/80 = $1.4

Price of one unit of carrots = 120/400 = $.3

Correct Answer:

100

The base year CPI is always 100.

Correct Answer:

Inflation rate in 2017 = 40%

Working note:

CPI in 2017 = (2.5*100 + 1.4*50+ .3*450)*100/(2.1*100+1.4*50+.1*450)

CPI in 2017 = 140

So, inflation rate = (140-100)/100 = 40%